nini01:

Janelle is in charge of bringing in dry erase markers for everyone in her math group to use to work on a project. She brings 3 markers for each person in her group and 5 extra markers for everyone to share. If Janelle brings in 23 markers in all, which equation represents the information? Let x represent the number of students in Janelle’s group.

crackhead17:

x is the number of students in her class

nini01:

3x = 23 + 5 3x = 23 3x – 5 = 23 3x + 5 = 23

xxDeppressionxx:

there are 6 students in Janelle's group. And x is 6 so do (3•6)+(5)=23

xxDeppressionxx:

So (D. .) would be you answer

nini01:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx So (D. .) would be you answer \(\color{#0cbb34}{\text{End of Quote}}\) thx

xxDeppressionxx:

Np